Currently, quarters have weights that are normally distributed with a mean of 5.670 grams and a standard deviation of 0.062 grams.

Question - Find the IQR of the weights of the quarters in that particular vending machine.
Answer #1

First quartile Q1 = 25th percentile

= Mean + Z * Standard deviation , Where Z is critical value at 25% confidence level.

= 5.67 + (-0.6745) * 0.062

= 5.628

Third quartile = 75th percentile

= Mean + Z * Standard deviation , Where Z is critical value at 75% confidence level.

= 5.67 + (0.6745) * 0.062

= 5.712

IQR = Q3 - Q1

= 5.712 - 5.628

= 0.084